OpenStudy (anonymous):

Olga is trying to figure out the length of a pool at a hotel that is to be built. A one inch door in the model will become a 3 foot door in reality. If the length of the pool is 7 inches on the model, how long will it be when it is built? 21 10 23 13

OpenStudy (alexrobin13):

So 1 inch = 3 feet. So if you have 7 inches how many inches is that in respect to that ratio?

OpenStudy (jdoe0001):

we know that for every 1 inch on the model equals 3 feet in the actual build 3 feet = 3 * 12 = 36 inches so \(\large { \begin{array}{ccllll} model&actual\ build \\\hline\\ 1&36\\ 7&x \end{array}\implies \cfrac{1}{7}=\cfrac{36}{x}\impliedby \textit{solve for "x"} }\)